OpenStudy (anonymous):

Solve for x in the first equation, then solve the system using substitution. Write your answer as (x, y) 2x-6y=14, 4x-y=9'

OpenStudy (anonymous):

Solve for x in the first equation, then solve the system using substitution. Write your answer as (x, y) 2x-6y=14, 4x-y=9'

OpenStudy (anonymous):

change second equation to y=4x-9 it would be easier

OpenStudy (anonymous):

then substitute that y into the first equation to get 2x-6(4x-9)=14 then distribute

OpenStudy (anonymous):

now you have 2x-24x+54=14 combine like terms to get -22x+54=14 isolate the variable by subtracting 54 from both sides now you have -22x=-40 divide both sides by -22 and your answer should be x=-40/-22=1.818181...

OpenStudy (anonymous):

(1.8181,1.7272)